OpenStudy (anonymous):

A spinner has three sections. The table shows the results of spinning the arrow on the spinner 80 times. Section 1 Section 2 Section 3 18 30 32 What is the experimental probability of the arrow stopping over Section 3?

OpenStudy (anonymous):

\[A. \frac{ 1 }{ 32 }\] \[B. \frac{ 3 }{ 32 }\] \[C. \frac{ 2 }{ 5 }\] \[D. \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

@inowalst

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Well, since the amount of times the spinner landed on Section 3 was 32 times, and the amount of times the spinner was spun is 80 times, the experimental probability of the spinner landing on Section 3 would most likely be 32/80, or simplified, it would be 2/5.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem. :)